OpenStudy (anonymous):

(5,3) (8,-2) in standard form?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

what does it mean "in standard form"? these are two coordinates

OpenStudy (anonymous):

i have to get them 2 point in a standard form as in Ax+Bx=C

OpenStudy (anonymous):

you mean you need the equation of the line through the two points yes?

OpenStudy (anonymous):

i guess i really dont no

OpenStudy (anonymous):

slope is \{\frac{-5}{3}\] because from 5 to 8 is three units to the right and from 3 to -2 is 5 units down. \[m=\frac{-2-3}{8-5}=\frac{-5}{3}\]

OpenStudy (anonymous):

maybe you need an equation f(x)=mx+b

OpenStudy (anonymous):

now use point -slope formula \[y-y_1=m(x-x_1)\] \[y-3=\frac{5}{3}(x-5)\]

OpenStudy (anonymous):

now multiply by 3 to get \[3y-9=-5(x-5)\] \[3y-9=-5x+25\] \[5x+3y=25\]